40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 14076     Entry time: Tue Jul 17 12:46:28 2018
Author: rana 
Type: Update 
Category: General 
Subject: some notes from yesterday 

For the EY, instead of balancing the table, I just moved the weight approximately so that the ETMY OSEMS were at half light, but didn't check the level since ETMY is the only optic.

 

Some notes on OMC/AS work (Aaron/Gautam can amend/correct):

- Beam is now well centered in OMC MMT. Hits input coupling mirror and cleanly exits the vacuum to the AS table.

- Didn't see much on OMC trans, but PDs are good based on flashlight test.

- just before closing, re-aligned beam in yaw so that it gets close to the east screw on the input coupler. Aaron and I think we maybe saw a flash there with the OMC length PZT being driven at full range by a triangle wave.

- with OMC Undulators (aka tip/tilt PZT mirrors) energized, the beam was low on PZT1 mirror. We pitched ITMY by ~150 micro-rad and that centered the beam on PZT1 mirror. ITMY-OL is probably not better than 100 urad as a DC reference?

- We checked the range of Undulator 1 and we were getting ~5 mrad of yaw of the beam for the full range, and perhaps half of that in pitch. Rob Ward emailed us from Oz to say that the range is robably 2.7 mrad, so that checks out.

Even if the ITMY has to be in the wrong position to get the beam to the OMC, we can still do the heater tests in one position and then do the OMC checkout stuff in the other position.

Gautam suspects that there is a possible hysterical behaviour in the Undulators which is related to the MC3 glitching and the slow machine hangups and also possibly the illuminati.

[aaron]

-We noticed a ghost beam that from MC REFL (MMT2) that should be dumped during the next vent--it travels parallel to the OMC's long axis and nearly hits one of the steering mirrors for OMC refl.

-We measured the level of the table and found it ~3 divisions off from level, with the south end tilted up

-Gautam rotated and slightly translated OM5 to realign the optic, as expected. No additional optics were added.

-Gautam and I tested the TT piezo driver. We found that 3.6V on the driver's input gave 75V (of 150V) at the output, at least for yaw on piezo 1. However, as Gautam mentioned, during testing it seemed that the other outputs may have different (nonzero) offset voltages, or some hysterisis.

ELOG V3.1.3-